OpenStudy (gabylovesyou):

Alex factored 32a^2 + 18b^2 using the following steps: 32a^2 + 18b^2 = 2(16a^2 + 9b^2) = 2(4a + 3b)(4a - 3b) Did he do it correctly? if not explain what he did wrong and how he can check his answers.

OpenStudy (anonymous):

He didnt do it correctly... 32a^2 + 18b^2 = 2(16a^2 + 9b^2) = 2(4a + 3b)(4a - 3b)*WRONG*! He forgot the negative sign in front of 3b... This is how he can check his answer 2(4a + 3b)(4a - 3b) = 2(16a^2-9b^2) = 32a^2-18b^2...which doesn't match the answer the answer they gave....

OpenStudy (gabylovesyou):

2(16a^2 + 9b^2) is correct but how did he get what goes next?

OpenStudy (anonymous):

Sorry its taking so long to respond...Im also helping others.

OpenStudy (anonymous):

One moment please.

OpenStudy (anonymous):

Yes that is correct...But what he got next is wrong...He used A2 - B2 = (A + B)(A - B)

OpenStudy (gabylovesyou):

ok

OpenStudy (gabylovesyou):

its supposed to be (a + b) (a - b) ?

OpenStudy (anonymous):

No thats what he used...And that was wrong..If you see when he times the +b and - b he should have had a -b^2 not a +b^2.

OpenStudy (gabylovesyou):

i sent u a PM... tell me if thats ok to put

OpenStudy (anonymous):

Lol yeah I got it, Its good! lol :)

OpenStudy (anonymous):

its impossible to factor out

OpenStudy (anonymous):

to show why, take last 2(4a + 3b)(4a - 3b) and try to remove brackets by multiplying each of the terms which i hope you can do yourself?

OpenStudy (anonymous):

@Tomas.A I already answered all her questions for this problem for her through messaging.

OpenStudy (anonymous):

ah thats ok, but she still was asking for help in math chat, no idea why, so just confirmed what you said

OpenStudy (anonymous):

Lol okay. :P Yeah I explained it and she got to the end, so I have no idea why. Thanks for letting me know! :)

OpenStudy (anonymous):

so you done with her expanding, right?

OpenStudy (anonymous):

Yes, I showed her everything and she even sent me a message making sure it was clear and correct...So she should be good. lol but I really dont know..You know how some people are just saying they get it when they dont and end up getting the answer without learning mostly anything. :P